OpenStudy (yanasidlinskiy):

If you know the number of yards for a measurement, you can change that measure to meters by multiplying the number of yards by 0.9144. a. Write a function rule that relates meters to yards. b. How many meters are equivalent to 7,200 yards? Round your answer to the nearest hundredth, if necessary. c. How many yards are equivalent to 2,500 meters? Round your answer to the nearest hundredth, if necessary.

hartnn (hartnn):

let yards be 'y' "multiplying the number of yards by 0.9144" meters = m = y * 0.9144

hartnn (hartnn):

that was your first part a for b. just plug in y =7200 in that equation m= y*0.9144

hartnn (hartnn):

for part c. since meters is given, plug in m = 2500 in m = y*0.9144 and find 'y'

OpenStudy (yanasidlinskiy):

Thanks soooo much!!:) That was helpful!!:)

hartnn (hartnn):

welcome ^_^